# “In Dialogues We Learn”: Towards Personalized Dialogue Without Pre-defined Profiles through In-Dialogue Learning

### v2

@@ 4.3 Evaluation Metrics
We employ various metrics to evaluate the performance of the dialogue models from the following aspects:
Coherence. BLEU-1/2 Papineni et al. (2002) and ROUGE-L Lin and Och (2004) are typical word overlap-based metrics for measuring the similarity between model responses and the ground-truth.
Diversity. Distinct-1/2 Li et al. (2015); Lv et al. (2023) consider the number of uni- or bi-grams in model responses, which are commonly used for evaluating diversity of dialogue generation.
Persona. Since our goal is to leverage persona information in dialogue sessions, we adopt P-F1 Ma et al. (2021) to measure the uni-gram F1 score between the model response and the latest utterance in the context. Inspired by Zhong et al. (2022), we use P-Co (Persona Cosine Similarity) as a supplement to the word overlap metrics to evaluate the semantic similarity between model responses and the ground-truth. Besides, following Tang et al. (2023), we also adopt Con.Score and Coh-Con.Score to measure the consistency between model responses and the given profiles in ConvAI2.
@@ 4.4 Main Results
CLV 11.85 15.1 71.24 92.89 71.72 23.01
LLaMA-2-7B IDL ${}^{\dagger}$${}^{\dagger}$${}^{\dagger}$ul96.97 ${}^{\dagger}$${}^{\dagger}$
LLaMA-2-7B gold 54.56 20.98 87.02 97.33 98.15 18.72
@@
LLaMA-2 IDL ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$
@@
LLaMA-2 IDL ${}^{\dagger}$ ${}^{\dagger}$ 10.62 ${}^{\dagger}$ ${}^{\dagger}$ 19.67 22.88
@@
LLaMA-2 IDL ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$ ${}^{\dagger}$${}^{\dagger}$
@@ 4.4.1 Automatic Evaluation
In Table 1, we compare the proposed method with existing personalized dialogue generation methods on ConvAI2. From the results, we can conclude that (1) when equipped with IDL, an open-source LLM can significantly outperform the existing methods in terms of almost all metrics, implying that IDL offers an effective way for leveraging LLMs in the task of personalized dialogue generation. (2) IDL can successfully recover personality characteristics from dialogue sessions. This is supported by the comparison between LLaMA-2 IDL and LLaMA-2 gold. Even without any hints from the profiles, IDL can still achieve comparable performance to the models fully supervised by the profiles.
@@ 4.4.2 Human Evaluation
We incorporate human evaluation to more accurately assess the quality of dialogues on three subjective dimensions: (1) Persona: evaluators will assess whether the response accurately and consistently reflects the persona information of the target person. (2) Style: evaluators will judge if the response aligns with the expected wording and tone for the target person. (3) Fluency: evaluators will examine the smoothness of the dialogue flow, considering both linguistic and logical fluency. We arranged the generated responses into pairs and conducted pairwise comparisons across these three dimensions.
Human evaluation results on ConvAI2 are shown in Figure 3. We sampled 500 pairs and engaged a professional evaluation group to perform the assessments. The two responses within each pair are produced from identical dialogue sessions and contexts, and the order of these two responses is randomized in the evaluation system. For each of the three dimensions mentioned previously, evaluators are required to assign a judgment of Win, Tie, or Lose based on the quality of these two responses.
The results show that IDL has brought significant improvements in both persona and style, with winning rates of 68.8% and 59.0% respectively, which demonstrates that the model using IDL can more effectively simulate the personality and tone of the target person. Regarding fluency, there is a slight decline in performance when using IDL, possibly attributed to the model’s increased focus on aligning with persona information.
@@ 4.5.1 Ablation Study
[…]
@@ Appendix A Appendix
A.1 Implementation Details
We experimented with a range of parameter combinations in our study. We adopt LLaMA-2-7B-Chat [^3] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf and LLaMA-2-13B-Chat [^4] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf as the backbones. The parameters utilized to obtain the experimental results presented in this chapter are as follows: In the MSL stage, the maximum number of clusters $c$ is set to $3$ and the maximum number of nearest neighbor $k$ is set to $5$. Scaling coefficient $\lambda$ is set to 5. We adopt Lora for training. The batch size is $4$ and the learning rate is $55e-$ 5. In the DPA stage, the penalty of DPOC is set to $2$. The batch size is set to 1 and the learning rate is $51e-$ 5. The model used as persona extractor is LLaMA-2-7B fine-tuned on PersonaExt. Our code is publicly available.  [^5] https://github.com/steven-ccq/In-Dialogue-Learning
A.2 convED
@@ A.1 Implementation Details
A.3 Case Study

### v4

@@ 4.3 Evaluation Metrics
We employ BLEU Papineni et al. (2002) and ROUGE-L Lin and Och (2004) metrics to assess the coherence of the text. [^5] We use NLTK to calculate both metrics. For evaluating diversity, Distinct-1/2 Li et al. (2015); Lv et al. (2023) metrics are utilized. Additionally, P-F1 Ma et al. (2021), P-Co (Persona Cosine Similarity) Zhong et al. (2022) are used to measure persona consistency, while Con.Score, and Coh-Con.Score are used to measure the consistency between model responses and the given profiles in ConvAI2 Tang et al. (2023).
@@ 4.4 Main Results
CLV 11.85 15.10 71.24 92.89 71.72 23.01
LLaMA2-7B System 7.22 9.56 72.21 94.39 98.87 22.32
[…]
@@
LLaMA-2 IDL 31.60 11.74 10.86 89.86 95.81 19.95 21.07
@@
LLaMA-2 IDL 32.56 13.00 10.62 90.31 97.24 19.67 22.88
@@
[…]
@@ 4.4.1 Automatic Evaluation
In Table 1, we compare the proposed method with existing personalized dialogue generation methods on ConvAI2. From the results, we can conclude that (1) when equipped with IDL, an open-source LLM can significantly outperform the existing methods in terms of almost all metrics, implying that IDL offers an effective way for leveraging LLMs in the task of personalized dialogue generation. (2) IDL can successfully discover persona information from dialogue sessions, comparing LLaMA-2 IDL with LLaMA-2 gold. Even without any hints from the profiles, IDL can still achieve comparable performance to the models fully supervised by the profiles.
@@ 4.4.2 Human Evaluation
We hire $5$ well-educated volunteers as annotators, and require them to judge the quality of model responses from three aspects: (1) Persona: the annotators assess whether a response accurately and consistently reflects the persona information of the target person. (2) Style: the annotators judge if the response aligns with the expected wording and tone for the target person. (3) Fluency: the annotators examine the smoothness of the dialogue flow, considering both linguistic and logical fluency. We sample $500$ dialogues associated with demonstrations from the test set of ConvAI2, and obtain responses for each dialogue from IDL and ICL (based on LLaMA-2-13B-Chat), respectively. Each time, a pair of responses are randomly shuffled and presented to the $5$ annotators. Each annotator assign labels from {Win, Tie, Lose} to a pair according to Persona, Style, and Fluency, and in total, each pair obtains $5$ labels on each of the three aspects. Figure 3 shows the evaluation results. IDL significantly improves upon ICL on both persona and style, with winning rates of 68.8% and 59.0%, respectively, demonstrating that the model using IDL can more effectively simulate the personality and tone of the target person. Regarding fluency, there is a slight decline in performance when using IDL, possibly attributed to the model’s increased focus on aligning with persona information. We calculate Cohen’s kappa, and the values for persona, style, and fluency are $0.53$, $0.56$ and $0.51$, respectively, indicating moderate agreement among the annotators.
@@ 4.5.1 Ablation Study
From the results, we observe that (1) DPOC plays a crucial role in enhancing the acquisition of better persona information, and the elimination of criterion samples significantly diminishes the model’s effectiveness. This is because the model can pay more attention to persona-related tokens after deep personalized alignment. Relevant case study can be found in Appendix A.5. Additionally, the findings suggest that merely employing DPO falls short in substantially improving the overall performance of models. This is because the preference alignment of DPO is not optimized for problems that can arise from personalized dialogue generation task, as illustrated in §$\S$ 3.3.2. Furthermore, the diminished effectiveness observed upon removing static and dynamic persona identifiers underscores the importance of reorganizing training data before the supervised fine-tuning process.
@@ 5 Conclusion
Acknowledgements
This work was supported by the National Natural Science Foundation of China (NSFC Grant No. U21B2027, 62122089), Beijing Outstanding Young Scientist Program NO. BJJWZYJH012019100020098, and Intelligent Social Governance Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China.
@@ Appendix A Appendix
A.1 Supplementary Results
We present the evaluation results in terms of BLEU-1/2/3/4 in Table 4. The prompt used in LLaMA2-System is “Here are your persona settings, your reply must be consistent with the persona: {profile}”. From the results, we can conclude that IDL holds consistent advantages over baseline methods on all BLEU metrics.
Dataset Size Model BLEU-1 BLEU-2 BLEU-3 BLEU-4
ConvAI2 124M MSP 8.19 2.34 1.56 0.73
125M CLV 11.85 4.29 2.41 1.02
7B LLaMA-2 IDL 52.40 25.43 12.74 8.40
7B LLaMA-2 gold 54.56 27.98 15.37 10.80
13B LLaMA-2 IDL 54.48 28.42 14.77 9.89
13B LLaMA-2 gold 55.32 28.71 16.00 11.33
Movie 7B Vicuna 14.76 5.53 2.36 1.29
7B LLaMA-2 ICL 6.12 3.07 1.36 0.72
7B LLaMA-2 IDL 31.60 11.74 4.89 2.87
13B Vicuna 12.82 4.01 1.43 0.70
13B WizardLM 29.60 10.45 6.03 3.83
13B LLaMA-2 ICL 15.04 7.00 3.65 2.07
13B LLaMA-2 IDL 32.56 13.00 5.56 3.32
LIGHT 7B Vicuna 36.07 17.37 7.42 3.83
7B LLaMA-2 ICL 15.41 8.92 3.99 2.02
7B LLaMA-2 IDL 46.32 22.01 9.63 5.27
13B Vicuna 19.68 8.87 3.53 1.75
13B WizardLM 44.59 21.45 10.30 5.53
13B LLaMA-2 ICL 24.31 13.47 6.00 3.08
13B LLaMA-2 IDL 49.69 24.64 10.90 6.03
Table 4: Evaluation results w.r.t. BLEU 1-4.
A.2 Implementation Details
In-Dialogue Learning.
In the Mutual Supervised Learning stage, the maximum cluster number $c$ is set to 3 and the maximum number of neighbors $k$ is set to 5. Besides, the scaling coefficient $\lambda$ is set to 5. We use LoRA for training. The rank is set to 8 and the lora_alpha is set to 8. We adopt AdamW as the optimizer. We set adam_beta1, adam_beta2 and adam_epsilon to 0.9, 0.999 and $1e^{-8}$, respectively. We use cosine schedule to warm up. The batch size is set to 4 and the learning rate is set to $5e^{-5}$. We finetune our model on ConvAI2 dataset for 2 epochs. Each epoch takes around 40 minutes. The training of this process is completed on one Nvidia A100 GPU.
In the Deep Personalized Alignment stage, we set the penalty of DPOC to 2. The batch size is set to 1 and the learning rate is set to $1e^{-5}$. We use LoRA for training. The rank is set to 8 and the lora_alpha is set to 8. We adopt AdamW as the optimizer. We set adam_beta1, adam_beta2 and adam_epsilon to 0.9, 0.999 and $1e^{-8}$, respectively. We use cosine schedule to warm up.
We fine-tune our models on the DPOC dataset, which is built based on ConvAI2. For each sample, we use the ground truth of ConvAI2 as the chosen sample, and select the one with the worst quality among the candidate responses (the candidate responses has been sorted by quality in the ConvAI2 provided by parlai) as the rejected sample. As for the criteria sample, we randomly select a type (line 386) and build it according to its corresponding construction method. The number of training epochs is set to 1. Each epoch takes around 12 hours. The training of this process is completed on one Nvidia A100 GPU.
Persona Extractor.
The original personaExt dataset contains 35K samples, with each sample containing a sentence and a triple <subject, relation, object>, which is a description of the persona information in the sentence. For example, the triple of the sentence "I have an apple, it is juicy." is <I, have, apple>. In order to adapt to the format of the conversation, we simply modified the original dataset by stacking multiple samples together to simulate the form of multiple sentences in a conversation. The modified dataset has 4K samples.
[…]
We use 90% of the data as the training set and the remaining 10% as the validation set. To test the accuracy of the trained Persona Extractor, for each sample in validation set, we concatenate sentences to form the input sequence. After that, we can get the output sequence of the Persona Extractor. We use the regex “<.*?,.*?,.*?>” to parse the output sequence. If one element of the triple is different from the ground truth, then the sample is judged “fals”. The accuracy of the Persona Extractor reached 87% in validation.
A.3 convED
@@ A.2 Implementation Details
A.4 Completed Loss Function for DPOC
Please refer to Equation 11.
A.5 Case Study
A.6 Low-resource Scenarios
We hope that the current conversation and its historical conversations are similar, so that the model can get more relevant information available from historical conversations. Therefore, in the Mutual Supervised Learning stage, we cluster input conversations so that they share more persona information (Static Persona Identification) and minimize the distance between the current conversation and historical conversations (Dynamic Persona Identification).
Of course, even so, we still can not guarantee that current conversation is similar to its historical conversations. Therefore, we add the following experiments. For each sample in original ConvAI2 test set, we replace the historical conversations with those in other clusters, so that the similarity between the historical conversations and the target conversation in this example is reduced. Results are shown in Table 5. The model used in this experiment is LLaMA-2 IDL 13B.
We can observe that the performance of the model has declined slightly. This is because the similarity between the historical conversations and the current conversation is the basic guarantee for IDL to be effective. When the current conversation involves a certain topic, IDL will focus on similar parts in historical conversations, thus completing the simulation of the persona. Therefore, when the similarity between the historical conversations and the current conversation decreases, the performance of IDL will also be affected.
A.7 Generalizability
To assess the generalizability of IDL, we also utilize LLaMA as the base model. Experimental results are presented in Table 6 and Table 7.
We repeated the training process of LLaMA-2 IDL using LLaMA and obtained LLaMA IDL. According to the results, LLaMA IDL surpasses Vicuna in multiple metrics, which further illustrates the effectiveness of IDL.
Size Similarity BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE-L Dist-1 Dist-2 Coh. Coh-Con
13B Original 54.48 28.42 14.77 9.89 20.05 87.78 97.45 98.48 19.63
13B Out-of-Cluster 52.37 25.93 11.69 6.97 16.49 90.84 98.95 97.82 6.60
7B Original 52.40 25.43 12.74 8.40 18.98 86.13 86.97 96.86 13.26
7B Out-of-Cluster 51.45 24.53 11.56 6.84 16.47 87.70 97.99 95.27 7.15
Table 5: Results for low-resource scenario.
$\mathcal{L}_{DPOC}(\pi_{\theta};\pi_{ref})=$ $-\mathbb{E}_{(x,y_{w},y_{r},y_{l})\sim\mathcal{D}}[\log\sigma(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{ref}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{ref}(y_{l}|x)})$ (11)
$-\min(0,\lambda\log\frac{\pi_{\theta(y_{w}|x)}}{\pi_{ref}(y_{w}|x)}-\lambda\log\frac{\pi_{\theta(y_{r}|x)}}{\pi_{ref}(y_{r}|x)})$
$-\min(0,\lambda\log\frac{\pi_{\theta(y_{r}|x)}}{\pi_{ref}(y_{r}|x)}-\lambda\log\frac{\pi_{\theta(y_{l}|x)}}{\pi_{ref}(y_{l}|x)})],$
@@ A.7 Generalizability
Model BLEU-1 BLEU-2 ROUGE-L Dist-1 Dist-2 P-F1 P-Co
Vicuna 7B 14.76 5.33 5.44 71.45 63.58 11.13 17.05
LLAMA 7B ICL 11.14 3.70 4.90 53.13 55.85 10.25 15.33
LLAMA 7B IDL 22.55 9.01 8.79 59.58 70.53 20.63 20.11
Vicuna 13B 12.82 4.01 3.88 75.37 60.53 6.54 14.22
LLAMA 13B ICL 11.27 3.74 4.42 45.85 46.49 8.83 14.85
LLAMA 13B IDL 24.11 9.69 8.79 68.02 78.43 18.25 20.91
Table 6: Generalizability experiment results on Movie dataset
Model BLEU-1 BLEU-2 ROUGE-L Dist-1 Dist-2 P-F1 P-Co
Vicuna 7B 36.07 17.37 10.52 83.27 90.56 16.53 23.4
LLAMA 7B ICL 19.39 7.80 6.83 61.36 64.83 10.75 17.01
LLAMA 7B IDL 46.89 23.18 13.87 80.68 93.48 24.07 31.21
Vicuna 13B 19.68 8.87 5.87 59.85 58.07 8.27 16.11
LLAMA 13B ICL 22.57 9.18 7.19 57.30 60.78 12.25 17.98
LLAMA 13B IDL 48.77 24.25 13.99 83.42 95.70 23.52 31.75
Table 7: Generalizability experiment results on LIGHT dataset